OpenStudy (anonymous):

Combine the like terms to simplify the expression: 9 + 3 + 17xy2 + 8y3 + 10x –13xy2 – 9x – 6y3 A. 2y3 + 4xy2 + x + 12 B. 2y6 + 4x2y2 + 10x + 12 C. 2y3 + 4xy2 + 10x + 9 D. 2y3 – 4xy2 – 3x + 12

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Like terms must have the same variables in the same powers.

Directrix (directrix):

9 + 3 + 17xy2 + 8y3 + 10x –13xy2 – 9x – 6y3 = 8y3 – 6y3 +17xy2 –13xy2 + 10x – 9x + 9 + 3 =

Directrix (directrix):

Look at the terms. The like terms are together. Be sure you see how they are grouped.

Directrix (directrix):

8y3 – 6y3 +17xy2 –13xy2 + 10x – 9x + 9 + 3 = ?

Directrix (directrix):

Waiting for you to add the like terms.

OpenStudy (anonymous):

I found out the answer it is C. 2y3 + 4xy2 + 10x + 9

OpenStudy (anonymous):

thx for your help

Directrix (directrix):

Let me see what I get.

Directrix (directrix):

I did not get that answer.

Directrix (directrix):

Whoever said this is wrong, I think. C. 2y3 + 4xy2 + 10x + 9

Directrix (directrix):

8y3 – 6y3 +17xy2 –13xy2 + 10x – 9x + 9 + 3 = 2y3 + 4xy2 + x + 12

Directrix (directrix):

@Mia_is_Awesome

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so it is A

OpenStudy (anonymous):

can you help me with another question

Directrix (directrix):

A looks okAy to me. Post the question in a new thread.

OpenStudy (anonymous):

ok